OpenStudy (anonymous):

\[z=3x^2\cos({\pi}xy)-3y^2\sin({\pi}x^2y)\\{x(t)=2t^2+1}\\{y(t)=3e^{2t}}\] calculate: \[{dz \over dt}\Bigr|_{t=0}\]

OpenStudy (anonymous):

What is the problem with this exercise?

OpenStudy (anonymous):

so to dz/dt means take the derivative of z. and for that you would have to use chain rule and product rule..

OpenStudy (anonymous):

ohh but it has to be t. ummm.

OpenStudy (anonymous):

my guess would be to plug x(t) and y(t) in the corresponding x and y of your z equation.

OpenStudy (zarkon):

\[\frac{dz}{dt}=\frac{\partial z}{\partial x}\frac{dx}{dt}+\frac{\partial z}{\partial y}\frac{dy}{dt}\]